cosh & sinh

what is cosh(npi) and sinh(npi) equal to??
is sinh(npi) =0 or cosh(0) =1
Answer #1

cosh(0) = e^0 + e^-0 /2 = 1+1/2 = 2/2 = 1

sinh(pi) = e^iPi - e^-ipi / 2 = 0 - 0 /2 = 0/2 = 0


answered by: xml